OpenStudy (anonymous):

Find the angle in radians between the planes -2x +z =1 and 3y+z =1. Really need to know how to do this.

OpenStudy (saifoo.khan):

@satellite73

OpenStudy (saifoo.khan):

Nope, sorry.

OpenStudy (anonymous):

Yeah I looked up the answer and it didn't work :/

OpenStudy (anonymous):

Do you know how to go about this it makes no sense to me

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Dihedral_angle I'm trying to work this out...

OpenStudy (anonymous):

Alright thanks a bunch!

OpenStudy (anonymous):

Here's a long shot...Since y doesn't occur in the first equation and x doesn't occur in the second one, does that mean that they are at 90° ?

OpenStudy (anonymous):

It has to be in radians and I have tried pi/2 and it doesnt seem to work :/

OpenStudy (anonymous):

http://mathworld.wolfram.com/DihedralAngle.html

OpenStudy (anonymous):

-2x +z =1 and 3y+z =1 n1 = -2,0,1 n2 = 0,3,1 n1.n2 = 1 cos(theta) = 1 theta = 0

OpenStudy (anonymous):

Oops, I forgot the denominator of the dot product...

OpenStudy (anonymous):

n1.n2 = 1/(sqrt(5)*sqrt(10)) = 1/(5sqrt(2)) arccos(1/(5sqrt(2)))

OpenStudy (anonymous):

You got it. You are my hero. Legit! Thanks a bunch!

OpenStudy (anonymous):

81.87 degrees?

OpenStudy (anonymous):

I just answer the arccos (1/(5sqrt(2))) and it was right haha

OpenStudy (anonymous):

I'm surprised it worked. I've never done those before.

OpenStudy (anonymous):

Haha well I am amazed thanks a lot!

OpenStudy (anonymous):

Thanks for the fun problem, see you!